# Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries

### v2

@@
Alexandros Skouras Institut Montpelliérain Alexander Grothendieck, Université de Montpellier, France alskouras@gmail.com Omar Lakkis Department of Mathematics, University of Sussex, Brighton UK lakkis.o.math@gmail.com  and  Vanessa Styles Department of Mathematics, University of Sussex, Brighton UK v.styles@sussex.ac.uk
(Date: March 18, 2025)
@@
We study a model for lithium (Li) electrodeposition on Li-metal electrodes that leads to dendritic pattern formation. The model comprises of a system of three coupled PDEs, taking the form of an Allen–Cahn equation, a Nernst–Planck equation and a Poisson equation. We prove existence of a weak solution and stability results for this system and present numerical simulations resulting from a finite element approximation of the system, which illustrate the dendritic nature of solutions to the model.
@@ 1. Introduction
In this paper we analyse a model for lithium (Li) electrodeposition on Li-metal electrodes in high-capacity lithium-oxygen and lithium-sulphur batteries. The model we consider is a variant of the thermodynamically consistent model introduced in Chen et al. (2015) to predict dendritic patterns which form during an electrochemical process using Li-electrodeposition. For ease of presentation we consider a simplified version in which, where possible, physical parameters have been set to $1$:
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.1)
$\partial_{t}c=\nabla\cdot\big{[}D(u)\nabla c+D_{1}(u,c)\nabla\phi\big{]}-\partial_{t}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.2)
$\nabla\cdot\big{[}\sigma(u)\nabla\phi\big{]}=\partial_{t}u-\sigma^{\prime}(u)\partial_{x}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.3)
@@
As humanity aims at reaching “net-zero” by 2050, replacing fossil fuels with renewable alternatives has spurred research in storage technologies, with lithium batteries being one of the favoured technologies. Electrodeposition in lithium batteries is an important phenomena that may lead to solid dendritic structures which, whilst remaining attached to the electrode, grow into the electrolyte solution. These structures can seriously compromise the performance and safety of such batteries, see for example Okajima et al. (2010); Akolkar (2013); Liang and Chen (2014); Chen et al. (2015); Yurkiv et al. (2018) and the references therein.
@@
In this paper we adapt techniques from Burman and Rappaz (2003) in which the authors establish an existence result for weak solutions to a simplified version of the model (1.1) - (1.3). Specifically, the model studied in Burman and Rappaz (2003), which does not include an equation of type (1.3), is a system of two equations that take similar forms to (1.1) and (1.2). There are two key differences between the forms of (1.1) and (1.2) in this paper and their respective forms in Burman and Rappaz (2003), firstly, in this work the $\partial_{t}u$ term in (1.2) is not included in Burman and Rappaz (2003) and secondly, the $\nabla\phi$ advection term in (1.2) in this work is replaced by a $\nabla u$ advection term in Burman and Rappaz (2003).
@@ 2.1. The model
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.1)
$\partial_{t}c=\nabla\cdot\big{[}D(u)\nabla c+D_{1}(u,c)\nabla\phi\big{]}-\partial_{t}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.2)
$\nabla\cdot\big{[}\sigma(u)\nabla\phi\big{]}=\partial_{t}u-\sigma^{\prime}(u)\partial_{x}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.3)
$\boldsymbol{{A}}(\nabla u)\nabla u\cdot\mathbf{n}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ D(u)\nabla c\cdot\mathbf{n}=D_{1}(u,c)\nabla\phi\cdot\mathbf{n}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ on }\partial\Omega\times[0,T],$ (2.4)
$\sigma(u)\nabla\phi\cdot\mathbf{n}=0\text{ on }\Gamma_{3}\times[0,T],\leavevmode\nobreak\ \phi=\phi_{-}\text{ on }\Gamma_{1}\times[0,T],\leavevmode\nobreak\ \phi=0\text{ on }\Gamma_{2}\times[0,T],$ (2.5)
$u(0)=u_{0},\leavevmode\nobreak\ \leavevmode\nobreak\ c(0)=c_{0}\text{ in }\Omega.$ (2.6)
@@
$\Gamma_{1}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=0\right\},\leavevmode\nobreak\ \Gamma_{2}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=L_{1}\right\}\text{ and }\Gamma_{3}=\partial\Omega\smallsetminus(\Gamma_{1}\cup\Gamma_{2}),$ (2.7)
$g(u):=u^{2}(1-u)^{2}/4,\leavevmode\nobreak\ h(u):=u^{3}(6u^{2}-15u+10),$ (2.8)
$D(u):=D^{\text{e}}h(u)+D^{\text{s}}(1-h(u)),\leavevmode\nobreak\ \leavevmode\nobreak\ \sigma(u)=\sigma^{\text{e}}h(u)+\sigma^{\text{s}}(1-h(u)),\leavevmode\nobreak\ \text{ for }u\in\mathbb{R},$ (2.9)
$m(c):=\left\{\begin{array}[]{ll}C_{1}-C_{2},&c>1,\\ C_{1}-cC_{2},&0\leq c\leq 1,\\ C_{1},&c<0,\end{array}\right.\leavevmode\nobreak\ \leavevmode\nobreak\ D_{1}(u,c):=\left\{\begin{array}[]{ll}D(u),&c>1,\\ D(u)c,&0\leq c\leq 1,\\ 0,&c<0,\end{array}\right.$ (2.10)
with $\operatorname{D}\!^{e},D^{s},\sigma^{e},\sigma^{s}\in\mathbb{R}$ and $0<D^{\text{s}}<D^{\text{e}}$, $0<\sigma^{\text{e}}<\sigma^{\text{s}}$. The anisotropic diffusion tensor-valued field $\boldsymbol{{A}}(\boldsymbol{p})$ of $\boldsymbol{p}\in\mathbb{R}^{d}\smallsetminus{\left\{{\boldsymbol{0}}\right\}}$ is defined in Section 2.4.
@@ 2.3. The weak formulation
$(\sigma(u)\nabla\phi,\nabla\eta)=(-\partial_{t}u+\sigma^{\prime}(u)\partial_{x}u,\eta),$ (2.11c)
@@
$g^{\prime}(s)=s(s-1)(s-\frac{1}{2}),\leavevmode\nobreak\ h^{\prime}(s)=30s^{2}(s-1)^{2}$ (2.12)
@@ 2.4. Anisotropic diffusion tensor
such that $J_{a}^{\prime}:\operatorname{H}^{1}(\Omega)\to(\operatorname{H}^{1}(\Omega))^{\prime},\leavevmode\nobreak\ \forall v\in\operatorname{H}^{1}(\Omega)$, satisfies
$J_{a}^{\prime}(w)[v]=\int_{\Omega}\boldsymbol{{A}}(\nabla w)\nabla w\cdot\nabla v,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall v\in\operatorname{H}^{1}(\Omega).$ (2.17)
@@ 2.5. Lemma (convexity of the anisotropic Dirichlet energy)
$\mathcal{H}(\boldsymbol{p})=\nabla^{2}\bar{a}^{2}(\boldsymbol{p})>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\boldsymbol{p}\in\operatorname{Dom}\bar{a}^{2},$
@@
With the use of software Wolfram Mathematica Wolfram Research, Inc. (2024), our calculations show that $\mathcal{H}_{11}$ is positive if and only if
@@
[…]
@@ 2.6. Remark (monotonicity)
$J_{a}(w)-J_{a}(v)\leq J_{a}^{\prime}(w)[w-v]=\int_{\Omega}\boldsymbol{{A}}(\nabla w)\nabla w\nabla(w-v).$ (2.24)
@@ 2.7. Remark (ellipticity)
$(\boldsymbol{{A}}(\boldsymbol{p})\boldsymbol{p},\boldsymbol{p})=(\nabla(\bar{a}^{2}(\boldsymbol{p})),\boldsymbol{p})=2\int_{\Omega}\bar{a}^{2}(\boldsymbol{p}),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\boldsymbol{p}\in\mathbb{R}^{d}$ (2.25)
thus from (2.15) and (2.19) we establish the bounds
$c_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\int_{\Omega}\boldsymbol{{A}}(\nabla w)|\nabla w|^{2}\leq C_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}$ (2.26)
@@
$c_{A}:=\min_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ C_{A}:=\max_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p}).$ (2.27)
@@ 3. Time discretization
In this section we consider a time discretization of (2.11a)–(2.11c). To this end, we partition the time interval $[0,T]$ into $K$ equidistant steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<t_{K}=T$, so that for $k=1,\ldots,K$, $t_{k}=k\tau$.
@@
$(d_{\tau}u^{k},\chi)+(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla\chi)+(g^{\prime}(u^{k}),\eta)=((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),\chi)$
$+((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi),$ (3.1a)
$(\sigma(u^{k})\nabla\phi^{k},\nabla\eta)=(-d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k},\eta),$ (3.1b)
@@
with $u^{0}=u(0)=u_{0}$, $c^{0}=c(0)=c_{0}$,
$h_{1}^{\prime}(u^{k}):=30u^{k-1}u^{k}(u^{k}-1)^{2}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ h_{2}^{\prime}(u^{k}):=30(u^{k})^{2}(u^{k}-1)(u^{k-1}-1)$ (3.2)
and $(\xi)^{-}:=\min(0,\xi)$ and $(\xi)^{+}:=\max(0,\xi)$.
@@
We make the following assumptions on the time step $\tau$, the anisotropy parameter $\delta$, defined in (2.15), and the initial data $u_{0},c_{0}$,

### v3

@@
Alexandros Skouras Institut Montpelliérain Alexander Grothendieck, Université de Montpellier, France alskouras@gmail.com , Omar Lakkis Department of Mathematics, University of Sussex, Brighton UK lakkis.o.math@gmail.com and Vanessa Styles Department of Mathematics, University of Sussex, Brighton UK v.styles@sussex.ac.uk
(Date: Ngày 17 tháng 12 năm 2025)
@@
We study a model for lithium (Li) electrodeposition on Li-metal electrodes that leads to dendritic pattern formation. The model comprises of a system of three coupled PDEs, taking the form of an Allen–Cahn equation, a Nernst–Planck equation and a Poisson equation. We prove existence of a weak solution and stability results for this system and present numerical simulations resulting from a finite element approximation of the system, which illustrate the dendritic nature of solutions to the model..
@@ 1. Introduction
As humanity aims at reaching “net-zero” by 2050, replacing fossil fuels with renewable alternatives has spurred research in storage technologies, with lithium batteries being one of the favoured technologies. Electrodeposition in lithium batteries is a phenomenon that plays an important role in studying lithium batteries, in that it may lead to solid dendritic structures which, whilst remaining attached to the electrode, grow into the electrolyte solution. Over time these structures can seriously compromise the performance, the lifespan as well as the safety of such batteries, see for example Okajima et al. (2010); Akolkar (2013); Liang and Chen (2014); Chen et al. (2015); Yurkiv et al. (2018) and the references therein.
In this paper we analyse a phase-field model for lithium (Li) electrodeposition on Li-metal electrodes in high-capacity lithium-oxygen and lithium-sulphur batteries. The model we consider is a variant of the thermodynamically consistent model introduced in Chen et al. (2015) to predict dendritic patterns which form during an electrochemical process using Li-electrodeposition. For ease of presentation we consider a simplified version in which, where possible, physical parameters have been set to $1$; namely, we look for functions $u,c,\phi$ of $(x,t)=(x_{1},\dotsc,x_{d},t)\in\varOmega\times T,d=2,3$ such that
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (1.1)
$\partial_{t}c=\nabla\cdot\big[D(u)\nabla c+D_{1}(u,c)\nabla\phi\big]-\partial_{t}u\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (1.2)
$\nabla\cdot\big[\sigma(u)\nabla\phi\big]=\partial_{t}u-\sigma^{\prime}(u)\partial_{x_{1}}u\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (1.3)
@@
In this paper we adapt techniques from Burman and Rappaz (2003) in which the authors establish an existence result for weak solutions to a simplified version of the model (1.1)–(1.3). Specifically, the model studied in Burman and Rappaz (2003), which does not include an equation of type (1.3), is a system of two equations that take similar forms to (1.1) and (1.2). There are two key differences between the forms of (1.1) and (1.2) in this paper and their respective forms in Burman and Rappaz (2003), firstly, in this work the $\partial_{t}u$ term in (1.2) is not included in Burman and Rappaz (2003) and secondly, the $\nabla\phi$ advection term in (1.2) in this work is replaced by a $\nabla u$ advection term in Burman and Rappaz (2003).
@@ 2.1. The model
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (2.1)
$\partial_{t}c=\nabla\cdot\big[D(u)\nabla c+D_{1}(u,c)\nabla\phi\big]-\partial_{t}u\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (2.2)
$\nabla\cdot\big[\sigma(u)\nabla\phi\big]=\partial_{t}u-\sigma^{\prime}(u)\partial_{x_{1}}u\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ in }\Omega\times[0,T],$ (2.3)
$\boldsymbol{{A}}(\nabla u)\nabla u\cdot\mathbf{n}=0,\penalty 10000\ \penalty 10000\ D(u)\nabla c\cdot\mathbf{n}=D_{1}(u,c)\nabla\phi\cdot\mathbf{n}\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ on }\partial\Omega\times[0,T],$ (2.4)
$\sigma(u)\nabla\phi\cdot\mathbf{n}=0\text{ on }\Gamma_{3}\times[0,T],\penalty 10000\ \phi=\phi_{-}\text{ on }\Gamma_{1}\times[0,T],\penalty 10000\ \phi=0\text{ on }\Gamma_{2}\times[0,T],$ (2.5)
$u(0)=u_{0},\penalty 10000\ \penalty 10000\ c(0)=c_{0}\text{ in }\Omega.$ (2.6)
@@
$\Gamma_{1}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=0\right\},\penalty 10000\ \Gamma_{2}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=L_{1}\right\}\text{ and }\Gamma_{3}=\partial\Omega\smallsetminus(\Gamma_{1}\cup\Gamma_{2}),$ (2.7)
$g(u):=u^{2}(1-u)^{2}/4,\penalty 10000\ h(u):=u^{3}(6u^{2}-15u+10),$ (2.8)
$D(u):=D^{\text{e}}h(u)+D^{\text{s}}(1-h(u)),\penalty 10000\ \penalty 10000\ \sigma(u)=\sigma^{\text{e}}h(u)+\sigma^{\text{s}}(1-h(u)),\penalty 10000\ \text{ for }u\in\mathbb{R},$ (2.9)
$m(c):=\left\{\begin{array}[]{ll}C_{1}-C_{2},&c>1,\\ C_{1}-cC_{2},&0\leq c\leq 1,\\ C_{1},&c<0,\end{array}\right.\penalty 10000\ \penalty 10000\ D_{1}(u,c):=\left\{\begin{array}[]{ll}D(u),&c>1,\\ D(u)c,&0\leq c\leq 1,\\ 0,&c<0,\end{array}\right$. (2.10)
with $\operatorname{D}\!^{e},D^{s},\sigma^{e},\sigma^{s}\in\mathbb{R}$ and $0<D^{\text{s}}<D^{\text{e}}$, $0<\sigma^{\text{e}}<\sigma^{\text{s}}$, and given constants $C_{1},C_{2}>0$. The anisotropic diffusion tensor-valued field $\boldsymbol{{A}}(\boldsymbol{p})$ of $\boldsymbol{p}\in\mathbb{R}^{d}\smallsetminus{\left\{{\boldsymbol{0}}\right\}}$ is defined in Section 2.4.
Note that the term $\sigma^{\prime}(u)\partial_{x_{1}}u$ in (2.3) appears from incorporating the boundary conditions for the potential $\phi$ in the equation. Keeping this term in (2.3) allows for simpler boundary conditions in the analysis, but in the numerical simulations, see Section 5, it will be added to the boundary conditions.
@@ 2.3. The weak formulation
$(\sigma(u)\nabla\phi,\nabla\eta)=(-\partial_{t}u+\sigma^{\prime}(u)\partial_{x_{1}}u,\eta),$ (2.11c)
@@
$g^{\prime}(s)=s(s-1)(s-\frac{1}{2}),\penalty 10000\ h^{\prime}(s)=30s^{2}(s-1)^{2}$ (2.12)
@@ 2.4. Anisotropic diffusion tensor
such that $J_{a}^{\prime}:\operatorname{H}^{1}(\Omega)\to(\operatorname{H}^{1}(\Omega))^{\prime},\penalty 10000\ \forall v\in\operatorname{H}^{1}(\Omega)$, satisfies
$J_{a}^{\prime}(w)[v]=\int_{\Omega}\boldsymbol{{A}}(\nabla w)\nabla w\cdot\nabla v,\penalty 10000\ \penalty 10000\ \forall v\in\operatorname{H}^{1}(\Omega).$ (2.17)
@@ 2.5. Lemma (convexity of the anisotropic Dirichlet energy)
$\mathcal{H}(\boldsymbol{p})=\nabla^{2}\bar{a}^{2}(\boldsymbol{p})>0,\penalty 10000\ \penalty 10000\ \forall\boldsymbol{p}\in\operatorname{Dom}\bar{a}^{2},$
@@
With the use of software Wolfram Mathematica (Wolfram Research, Inc., 2024), our calculations show that $\mathcal{H}_{11}$ is positive if and only if
@@
$1-15\delta>0,\penalty 10000\ \penalty 10000\ 1+6\delta-315\delta^{2}>0,\penalty 10000\ \penalty 10000\ 1+18\delta-15\delta^{2}>0,$ (2.22)
@@ 2.6. Remark (monotonicity)
$J_{a}(w)-J_{a}(v)\leq J_{a}^{\prime}(w)[w-v]=\int_{\Omega}\nabla w\cdot\boldsymbol{{A}}(\nabla w)\nabla(w-v).$ (2.24)
@@ 2.7. Remark (ellipticity)
$(\boldsymbol{{A}}(\boldsymbol{p})\boldsymbol{p},\boldsymbol{p})=(\nabla(\bar{a}^{2}(\boldsymbol{p})),\boldsymbol{p})=2\int_{\Omega}\bar{a}^{2}(\boldsymbol{p}),\penalty 10000\ \penalty 10000\ \forall\boldsymbol{p}\in\mathbb{R}^{d}$ (2.25)
thus from (2.15) and (2.19) we derive the bounds
$c_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\int_{\Omega}\left|\boldsymbol{{A}}(\nabla w)\right||\nabla w|^{2}\leq C_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}$ (2.26)
@@
$c_{A}:=\min_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p})\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ C_{A}:=\max_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p}).$ (2.27)
@@ 3. Time discretization
In this section we study a time discretization of (2.11a)–(2.11c). To this end, we partition the time interval $[0,T]$ into $K$ equidistant steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<t_{K}=T$, so that for $k=1,\ldots,K$, $t_{k}=k\tau$.
@@
$(d_{\tau}u^{k},\chi)+(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla\chi)+(g^{\prime}(u^{k}),\eta)=((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),\chi)$
$+((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),\chi),$ (3.1a)
$(\sigma(u^{k})\nabla\phi^{k},\nabla\eta)=(-d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x_{1}}u^{k},\eta),$ (3.1b)
@@
with $u^{0}=u(0)=u_{0}$, $c^{0}=c(0)=c_{0}$, $(\xi)^{-}:=\min(0,\xi)$, $(\xi)^{+}:=\max(0,\xi)$,
${\textsl{H}}_{1}(u^{k},u^{k-1}):=30u^{k-1}u^{k}(u^{k}-1)^{2},\penalty 10000\ \penalty 10000\ {\textsl{H}}_{2}(u^{k},u^{k-1}):=30(u^{k})^{2}(u^{k}-1)(u^{k-1}-1).$ (3.2)
3.1. Remark (Right-hand side of (3.1))
We can retrieve $h^{\prime}(s)$ as defined in (2.12) by setting $r=s$ in $\textsl{H}_{1}(s,r)$ and $\textsl{H}_{2}(s,r)$ such that
$h^{\prime}(s)=\frac{1}{2}\big(\textsl{H}_{1}(s,s)+\textsl{H}_{2}(s,s)\big).$ (3.3)
Associated with $\textsl{H}_{1}(s,r)$ and $\textsl{H}_{2}(s,r)$ are two functions $h_{1},h_{2}:\mathbb{R}^{2}\to\mathbb{R}$, which are antiderivatives in $s$ of $\textsl{H}_{1}$ and $\textsl{H}_{2}$, such that
$\begin{gathered}h_{1}(s,r):=15r\Big(\frac{(s-1)^{4}}{4}+\frac{(s-1)^{3}}{3}\Big)+\frac{5r}{4}\\ h_{2}(s,r):=15(r-1)\Big(\frac{s^{4}}{4}+\frac{s^{3}}{3}\Big)-\frac{5r}{4}\end{gathered}$ (3.4)
where, noting (2.8), we have
$h(s)=h_{1}(s,s)+h_{2}(s,s).$ (3.5)
@@
We make the following assumptions on the time step $\tau$, anisotropy parameter $\delta$, defined in (2.15), and the initial data $u_{0},c_{0}$,